The sum of the digits of a two two-digit number is 13. If the digits are reversed, the new number is 27 less than the original number. Find the number.
Choices:
76
94
95
85

Answer by mananth(16949)

x in the tens place
y in the digits place
the number be 10x+y
x+y =13

reversed number = 10y+x
10y+x = 10x+y -27
9y-9x=-27
/9
y-x=-3
add the two equations
2y = 10
/2
y=5
x+y=13
y=5
x=8
85 is the number